OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. -1/16x^2=y

OpenStudy (anonymous):

lol everyone hates these conic section questions don't they ? start maybe with \[x^2=-16y\] which looks a lot like \[(x-h)^2=4p(y=k)\] with \(h=k=0\) meaning the vertex is at the origin \((0,0)\) and \(p=-4\)

OpenStudy (anonymous):

actually i had a typo there, should have been \((x-h)^2=4p(y-k)\)

OpenStudy (anonymous):

i have this so far: but i think im wrong, Vertex: (0, 0); Focus: (-8, 0)

OpenStudy (anonymous):

think the focus is wrong \(p=-4\) so go 4 units down from \((0,0)\) to get the focus

OpenStudy (anonymous):

Focus: (0, -4)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and directix is the horizontal line 4 units up

OpenStudy (anonymous):

Vertex: (0, 0); Focus: (0, -4); Directrix: y = 4; Focal width: 16

OpenStudy (anonymous):

could be, i have no idea what "focal width" is, but i could look it up

OpenStudy (anonymous):

hahha is eveyrhting else i did right?

OpenStudy (anonymous):

yes, and focal width is \(|4p|=16\) so you are right there as well

OpenStudy (anonymous):

THANK YOU